Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 556
Entry time: 14 Oct 2024, 09:41
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 14.10.2024 |
|
|
General Announcement:
- Currently machine is now operating at 16.3GeV.
- The Winter maintenance task list was published last week. In case of questions please reach XO and/or Martin Doering.
CTRL:
- Issue on Thursday causing for the need to restart broker topic at MID and HED: caused by maintenance activity and triggered (unlikely) a chain of circumstances. Now it is fixed and understood.
The fix still has to be deployed and is unlikely that will happen again during in operation delivery.
DA:
- Nothing to report.
EEE:
- Couple of calls from PLC colleagues, but nothing major to report.
VAC:
- Nothing to report.
XPD:
- Nothing to report.
DET:
- Nothing major to report; preparing for tests of the detector in cw45 at HED - still to be scheduled.
ITDM:
- Maintenance on Dcache (only old data) today between 10:00 and 12:00.
XRO:
- Nothing to report.
LAS:
- Nothing to report.
UNSYS:
- Nothing to report.
SPB:
- Nothing to report.
FXE:
- Preparing for the night shifts; need higher pulse intensity from PP-laser that last week.
MID:
- Some devices are still down due to the broker issue last week. One has to figure out a bit more in the course of the day. CTRL will help.
HED:
- Nothing to report.
SCS:
- Not present.
SQS:
- Nothing to report.
SXP:
- Not present. |